Preform the indicated operation. Wright the answer in scientific notation.
( 6 x 10^-2)(5.67 x 10^3)

This is how I put this into the calculator and I got 385.344. I move the decimal over two places to get the answer 3.85 x 10^2. What happened to the .344?
I am studying for a test and I have the answers for the pre-test and I am allowed to use the calculator on the test. I t will help so much if I really knew how to use it.

Answer by Alan3354(69443)

( 6 x 10^-2)(5.67 x 10^3)
= 6*5.67*10^1
= 34.02*10
= 3.402*10^2
-------------
How did you get 385.344 ?